OpenStudy (anonymous):

You are driving from town A to town B. From town A you go 3 miles West and 4 miles South. What is the distance between the two towns? 1) 7 2) 5 3) 1 4) 2

OpenStudy (anonymous):

 Find d.

OpenStudy (saifoo.khan):

d is displacement. :D

OpenStudy (anonymous):

Just use the Pythagorean Theorem. \[c^{2} = a^{2} + b^{2}\]

OpenStudy (saifoo.khan):

@Calcmathlete , so distance or displacement? :D

OpenStudy (anonymous):

lol. Distance? :D

OpenStudy (saifoo.khan):

distance = 7 displacement = 5

OpenStudy (saifoo.khan):

lol. i think the teacher gave a wrong question.

OpenStudy (anonymous):

oh. I had a brain fart moment ;)

OpenStudy (saifoo.khan):

Haha.

OpenStudy (anonymous):

Are you two done having fun, and yes i got the joke. Haha... very funny P.S: Math is not strong subject

OpenStudy (anonymous):

Well, that kinda depends on if you figured it out yet? We're waiting to see you through until you ge the correct answer :) c^2 = a^2 + b^2 c^2 = 3^2 + 4^2